OpenStudy (anonymous):

A=(1/2)bh Solve the formula for h. Please show the steps of how you got the answer!

OpenStudy (anonymous):

A=frac{b*h}{2} \[2A=b*h\]

OpenStudy (anonymous):

\[A=\frac{b*h}{2} \]

OpenStudy (anonymous):

h=2A/b

OpenStudy (luffingsails):

Hi Epaema, In order to solve for h, we concentrate our efforts on what is on the same side as h. Our effort is to get h all alone, right? So, we start with: \[A = (1/2)bh\] So, start by multiplying both sides by 2 (what we do to one side has to be done to the other). We are multiplying by 2 to get rid of the 1/2 on the right. \[2A = 2(1/2)bh\] Which simplifies to: \[2A=bh\] Now, just divide each side by b and h will be all alone. \[(2A)/b=(bh)/b\] Simplifying a final time: \[h = 2A/b\] All done!

OpenStudy (anonymous):

THANK YOU BOTH! Even though luffingsails response was a lot easier to understand, you both got the same end result, so thank you! :)